### v1

@@ 4.2 Magnetic field orientation versus density structure of Oph-A
To investigate what is the relative orientation of the B-fields with respect to the gas structure in Oph-A, we quantify the projected offset angle between the orientations of B-field vectors projected in the plane-of-the-sky and the main structure associated with the ridge of the cloud, given as , where is the position angle of the ridge. We identify the shape of the ridge using the Python package RadFil444https://github.com/catherinezucker/radfil (Zucker & Chen, 2018, see appendix F).
@@
Figure 9 shows the map of offset angles () in Oph-A, overlaid by contours of continuum emission at 154 m. Clearly, the offset angle appears to increase when moving from the lower-density regions to the denser regions, indicating that the relative orientation of B-fields with respect to the main structure of the cloud turns from parallel to perpendicular toward the high-density region. This is in agreement with Planck polarization observations at 353 GHz (850 m) toward the entire L1688 region where the relative orientation of B-fields with respect to the gas structure changes gradually from parallel in the low-density region to perpendicular toward the higher-density region, with the transition occurring at cm-2 corresponding to the visual extinction of 3.3 mag555We convert from using the relation: molecules cm-2 mag-1 with a “standard” assumption of the total-to-selective extinction ratio, , of 3.1 (Frerking et al., 1982). (Soler, 2019). This transition is also observed in other nearby molecular clouds in the Gould Belt using Planck 353 GHz polarization observations, typically occurring at cm-2 or 2.7 mag (see e.g., Planck Collaboration et al., 2016; Soler et al., 2017). This phenomenon can be explained by the motion of the high-density material during the self-gravitational collapse. When the contraction occurs, the material becomes denser and compressed, which could cause the B-field lines to be re-configured. Eventually, the B-field lines are dragged and reshaped to become more uniform and perpendicular to the high-density structure of the field.
@@
The HAWC data traces B-fields in the dense region of Oph-A with cm-2 (see Figure 1), which corresponds to in range of 3.2–256 mag. Thus, we further investigate in detail the relation between the relative orientation of B-fields with respect to the cloud structure and the within Oph-A, as shown in Figure 10. Clearly, in the regime where 30 mag the offset angle () decreases with and drops from 60° at =4 mag down to 0° at 20 mag. This indicates that the relative orientation of B-fields changes from perpendicular back to parallel at 18–33 mag. Similar trends of this transition were found in the dense region of the Serpens South filament (Pillai et al., 2020) and in Serpens Main dense cloud (Kwon et al., 2022), occurring at 21 mag and 50 mag, respectively. This is likely because the B-fields are already dragged into the flow of very dense material during the gravitational contraction. Interestingly, we observe that the offset angle then progressively increases with 40 mag in our region, indicating the relative orientation of B-field with respect to the ridge changes back from parallel to perpendicular (with the transition at 42–60 mag). We note, however, that this feature was not seen in the Serpens South filamentary cloud whose maximum 60 mag (see Figure 3 in Pillai et al. (2020)) but in the Serpens Main cloud region where 170 mag (see Figure 6 in Kwon et al. (2022)).
@@ 4.3 Magnetic field versus gas density
The B-field strength ( or ) can be expressed as a function of (): . Theoretical models predict 2/3 in the weak B-field regions (Mestel, 1966) and 0.5 in the strong B-field regions where magnetic energy dominates over the gravity (Mouschovias & Ciolek, 1999). Consequently, in the latter case, it suggests that would increase faster than if there is any collapse in the dense cloud.
Here, we show in Figure 11 the relation toward the Oph-A cloud, using B-fields measured with the DCF method in band D. We find that for densities , the B-fields distribution tends to be flattened. In a regime with densities higher than we start seeing the increasing trend of B-fields with respect to gas densities. Thus, we decide to test the relation at the densest part of the cloud only. We choose this region as the area within the third contour of Stokes- at band D (see in the bottom right panel of Figure 5), where Stokes. Within this region, it reveals a relatively weak correlation between the B-fields and densities (with the Pearson coefficient of the correlation is found , corresponding to a significance of ). We perform the power-law fit using the Python curve-fit routine (), using mean values of estimated within every interval of =0.08666The bin width is chosen by using the Freedman-Diaconis rule. , which results in a power-law index of . In Figure 11, we also overplot the results of using other methods in subregions (a) and (e). We find that the fitted line well covers the data point representing the resulting value using ADF method and shows good agreement with those estimated with ST and DMA in subregion (a) within of uncertainty. The resulting is found consistent with theoretical predictions for strong B-fields (Mouschovias & Ciolek, 1999). This is also consistent with , representing strong B-fields in Oph-A (see bottom panel in Figure 8 and Section 3.6). We further perform the fit to values measured with the ST method from Sect. 3.3.2 and find . We find a similar power-law index for the dataset at band C, which are and using result of and , respectively (see Appendix E).
@@ 4.4 Magnetic fields versus stellar feedback
High-mass stars irradiate, heat up, and ionize the surrounding gas and create the H ii region (see e.g., Zinnecker & Yorke, 2007). The ionized, hot gas also expands and interacts with the vicinity environment. The radiation pressure acting on both dust and gas (assuming dust is well-coupled to the gas in the cloud) can drive the material away into the denser region. The lower-density matter in the outer regions of Oph-A is more exposed to the ionization front and pushed away faster than the higher-density matter in the inner regions. This results in dust and gas being compressed and forming the dense shell structure at the edge of the cloud. Here, we investigate a possible impact of the radiation field from the high-mass star Oph-S1 on the Oph-A cloud. The Oph-A molecular cloud is directly exposed to the strong UV radiation field emitted by the Oph-S1 star from the East. Figure 12 shows the map of the continuum emission at 154 m from SOFIA/HAWC+, superimposed by the white contours showing the H2 column density.
@@
We calculate the external radiation pressure of the Oph-S1 star acting on dust and gas in the densest part of the cloud, given by:
@@
where =1100 is the luminosity of the Oph-S1 star adopted from Lada & Wilking (1984) and 0.06 pc is the projected distance to the Oph-S1 star. We obtain K cm-3, which is in agreement with the result calculated from the ionizing photons emitted by the Oph-S1 star (2 K cm-3, Pattle et al., 2015). Here, we omit the fact that the UV radiation from the Oph-S1 star can be absorbed by intervening dust in between the star and the Oph-A region. Thus, the actual acting on dust and gas in Oph-A is overestimated in the Eq. 16. To compare with the external radiation pressure, we calculate the internal thermal pressure due to the hot gas in Oph-A, . We consider the hot gas within the densest part of Oph-A where ¿0.4 Jy/arcsec2 at 154 m only, which corresponds to the area inside the third contour shown in bottom right panel of Figure 5. Within this region, we adopt in the range of 104–106 cm-3 and the average temperature of the hot gas 20 K, which yields in the range of (0.2–20) K cm-3. With this assumption, we note that could be underestimated since there is a gas heating process in the cloud dominated by the UV radiation from the Oph-S1 star and by the photoelectric emission from dust exposed to the UV radiation (i.e., ). The highest internal thermal pressure of 20106 K cm-3 is seen toward the densest part of the cloud, associated with the position of the SM1 core. This suggests that the internal thermal pressure is strong enough to help the material refrain from external pressure. We note that, however, in the vicinity region away from the center (for e.g., the tails of the ridge shown in Figure 12), the internal thermal pressure is weaker, at least one order of magnitude less than the external one. In this case, the radiation pressure from the Oph-S1 star on the east side is dominant and able to push away all the matter into the surrounding environment toward the west side. However, it is clear that the material in the low-density region still survives, despite owning a low internal thermal pressure (see Figure 12), indicating that there must be another mechanism at play, for example, the presence of strong B-fields of a few mG in the cloud. We estimate roughly the magnetic pressure () to be 107–108 K cm-3 for B-field strengths between 1–2 mG. This indicates that the magnetic pressure is at least two orders of magnitude higher than the external pressure of 106 K cm-3, sufficient to support the matter against the external pressure along with the gas thermal pressure.
@@
We further investigate the interplay between B-fields and radiation directions on the evolution of the cloud. We determine the main radiation direction from the Oph-S1 star toward the densest region associated with the position of the starless core SM1 as the sky-projected direction from the Oph-S1 star and SM1 core. We find that the main radiation direction is 70°, measured east of north. The main B-field orientation toward the densest region of Oph-A of 74°, almost parallel to the radiation field. The strong radiation pressure from the Oph-S1 star pushes all the material away and creates a dense shell toward the west side. To prevent dust and gas compression in the dense region, the B-fields exert a magnetic pressure perpendicular to the B-field orientation, seen perpendicular to the direction of the radiation pressure. This suggests that the magnetic pressure does not help the matter resist the external pressure, but tends to stretch the structure out along the northwest-southeast direction, resulting in an elongated shape of the dense shell. In the west side, the external thermal pressure is blocked by the gas thermal pressure due to the high-density material, resulting in the ridge shape in the cloud with very high densities. These results are consistent with results from 3D MHD simulation, which predicted that if the B-field orientation is parallel or along with the radiation direction, the cloud will tend to have a flattened shape Henney et al. (2009). However, in the northeast and southeast regions the offset angles between the B-field vectors and radiation field are larger, i.e., the B-field orientation becomes perpendicular to the radiation direction. In this case, the magnetic pressure and tension induced by B-fields run parallel with the radiation pressure, suggesting magnetic pressure can support the matter to confront the radiative compression and prevent the sweeping dust and gas away. This results in bending the material structure and forming a dense ridge with an umbrella-like shape. Interestingly, in the case study of Oph-A the ionization source is very close to the cloud (¡0.1 pc) and the offset position angle between the radiation and B-fields is not homogeneous but varies from parallel to orthogonal. This allows us to examine widely how the interplay between the B-fields and radiation fields could influence the material structure of the cloud.
@@ 5 Summary
In this paper, we presented the analysis using the SOFIA/HAWC polarimetric observations toward the Oph-A molecular cloud in the L1688 active star-forming region of the Ophiuchus dark cloud complex. Our main results are as follows:
@@
Using the dust polarization maps at 89 and 154 m, we derived the map of B-field morphology in Oph-A. We found that B-fields are generally well-ordered and primarily perpendicular to the ridge of the cloud toward the higher-density regime. We estimated in the entire region of Oph-A using the DCF method, ranging from 0.2 to 2.3 mG at 89 and to 2.5 mG at 154 m. We found is prominent toward the densest region of the cloud associated with the starless core SM1 and decreases toward the outskirts. This is the first time we estimate the distribution of the B-field strengths in Oph-A, which is essential to understanding how B-fields vary in the molecular cloud globally. Our derived B-fields morphology and strengths well agree with previous measurements using the sub-mm dust polarization (Kwon et al., 2018). We also estimated strength using more advanced techniques, including the ADF, ST, and DMA, toward two subregions associated with the densest part of the cloud. While strengths estimated with the ADF method were in agreement with those from DCF, the ST and DMA methods provided lower values of 4.5 and 2-3 times, respectively.
@@
We studied the role of B-fields relative to gravity and turbulence by calculating the maps of the mass-to-flux ratio, Alfvénic Mach number, and plasma parameter toward Oph-A using the inferred from SOFIA/HAWC data at 154 m. We found that toward the central dense parts the cloud is magnetically sub-critical and sub-Alfvénic where B-fields dominate over the gravity and turbulence. We obtained in Oph-A, suggesting the cloud is supported by the strong magnetic pressure and much greater than the thermal gas pressure. The lower strengths obtained with the ST and DMA methods do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation. Our virial analysis in the entire cloud showed the virial parameter 15, suggesting that the cloud is gravitationally unbound and able to against the self-gravitational collapse. The cloud may be in the magneto hydrostatic equilibrium and may undergo gravitational collapse due to turbulent fluctuations on smaller scales. We should notice that our rough estimation of is a global parameter, applied for the entire cloud. In contrast, Oph-A is known as one of the densest parts of the active star-forming region L1688 hosting a few pre- and proto-stellar cores and embedded protostars. Higher spatial resolution observations toward the smaller scale region in the cloud will be necessary to investigate in detail the dynamical state of the cloud.
@@
We examined the role of B-fields in supporting the cloud against the radiation pressure from the nearby high-mass star Oph-S1. We found that at the center part of the cloud, the magnetic pressure indeed does not play a role in helping the material prevent radiation pressure. However, toward the northeast and southeast regions of the cloud, the magnetic pressure positively supports the cloud to prevent the radiative pressure together with the gas thermal pressure.

### v2

@@ 4.2 Magnetic field orientation versus density structure of Oph-A
To investigate what is the relative orientation of the B-fields with respect to the gas structure in Oph-A, we quantify the projected offset angle between the orientations of B-field vectors projected in the plane-of-the-sky and the main structure associated with the ridge of the cloud, given as , where is the position angle of the ridge. We identified the shape of the ridge using the Python package RadFil444https://github.com/catherinezucker/radfil (Zucker & Chen, 2018, see Appendix E).
@@
Figure 9 shows the map of offset angles () in Oph-A, overlaid by contours of continuum emission at 154 m. Clearly, the offset angle appears to increase when moving from the lower density regions to the denser regions, indicating that the relative orientation of B-fields with respect to the main structure of the cloud turns from parallel to perpendicular toward the high-density region. This is in agreement with Planck polarization observations at 353 GHz (850 m) toward the entire L1688 region where the relative orientation of B-fields, with respect to the gas structure changes gradually. It runs parallel in the low-density region to perpendicular toward the higher density region, with the transition occurring at cm-2 and corresponding to the visual extinction of 3.3 mag555We convert from using the relation: molecules cm-2 mag-1 with a “standard” assumption of the total-to-selective extinction ratio, , of 3.1 (Frerking et al., 1982). (Soler, 2019). This transition has also been observed in other nearby molecular clouds in the Gould Belt using Planck 353 GHz polarization observations, typically occurring at cm-2 or 2.7 mag (see e.g., Planck Collaboration et al., 2016; Soler et al., 2017). This phenomenon can be explained by the motion of the high-density material during the self-gravitational collapse. When the contraction occurs, the material becomes denser and compressed, which could cause the B-field lines to be re-configured. Eventually, the B-field lines are dragged and reshaped to become more uniform and perpendicular to the high-density structure of the field.
@@
The HAWC data traces B-fields in the dense region of Oph-A with cm-2 (see Fig. 1), which corresponds to in range of 3.2–256 mag. Thus, we further investigate in detail the relation between the relative orientation of B-fields with respect to the cloud structure and the within Oph-A, as shown in Fig. 10. Clearly, in the regime where 30 mag the offset angle () decreases with and drops from 60° at =4 mag down to 0° at 20 mag. This indicates that the relative orientation of B-fields changes from perpendicular back to parallel at 18–33 mag. Similar trends of this transition were found in the dense region of the Serpens South filament (Pillai et al., 2020) and in the Serpens Main dense cloud (Kwon et al., 2022), occurring at 21 mag and 50 mag, respectively. This is likely because the B-fields are already dragged into the flow of very dense material during the gravitational contraction. Interestingly, we observe that the offset angle then progressively increases with 40 mag in our region, indicating the relative orientation of B-field with respect to the ridge changes back from parallel to perpendicular (with the transition at 42–60 mag). We note, however, that this feature was not seen in the Serpens South filamentary cloud, whose maximum is 60 mag (see Fig. 3 in Pillai et al. (2020)); however, it was seen in the Serpens Main cloud region, where 170 mag (see Fig. 6 in Kwon et al. (2022)).
@@ 4.3 Magnetic field versus gas density
The B-field strength ( or ) can be expressed as a function of (): . Theoretical models predict 2/3 in the weak B-field regions (Mestel, 1966) and 0.5 in the strong B-field regions, where magnetic energy dominates over the gravity (Mouschovias & Ciolek, 1999). Consequently, in the latter case, it suggests that would increase faster than if there is any collapse in the dense cloud.
Here, we show in Fig. 11 the relation toward the Oph-A cloud, using B-fields measured with the DCF method in band D. We find that for densities , the B-fields distribution tends to be flattened. In a regime with densities higher than we start seeing the increasing trend of B-fields with respect to gas densities. Thus, we decided to test the relation at the densest part of the cloud only. We chose this region as the area within the third contour of Stokes- at band D (see in the bottom right panel of Fig. 5), where Stokes. Within this region, it reveals a relatively weak correlation between the B-fields and densities (with the Pearson coefficient of the correlation is found , corresponding to a significance of ). We performed the power-law fit using the Python curve-fit routine (), using mean values of estimated within every interval of =0.08666The bin width is chosen by using the Freedman-Diaconis rule. . This gives a power-law index of . In Fig. 11, we also overplot the results of using other methods in subregions (a) and (e). We find that the fitted line well covers the data point representing the resulting value using ADF method and shows good agreement with those estimated with ST and DMA in subregion (a) within and uncertainty of . The resulting is found consistent with theoretical predictions for strong B-fields (Mouschovias & Ciolek, 1999). This is also consistent with , representing strong B-fields in Oph-A (see bottom panel in Fig. 8 and Sect. 3.6). We then performed the fit to values measured with the ST method from Sect. 3.3.2 and find . We found a similar power-law index for the dataset at band C, namely, and using result of and , respectively (see Appendix C).
@@ 4.4 Magnetic fields versus stellar feedback
High-mass stars irradiate, heat up, and ionize the surrounding gas and create the H ii region (see e.g., Zinnecker & Yorke, 2007). The ionized, hot gas also expands and interacts with the vicinity environment. The radiation pressure acting on both dust and gas (assuming dust is well-coupled to the gas in the cloud) can drive the material away into the denser region. The lower density matter in the outer regions of Oph-A is more exposed to the ionization front and pushed away faster than the higher density matter in the inner regions. This results in dust and gas being compressed and forming the dense shell structure at the edge of the cloud. Here, we investigate a possible impact of the radiation field from the high-mass star Oph-S1 on the Oph-A cloud. The Oph-A molecular cloud is directly exposed to the strong UV radiation field emitted by the Oph-S1 star from the east. Figure 12 shows the map of the continuum emission at 154 m from SOFIA/HAWC+, superimposed by the white contours showing the H2 column density.
@@
We calculated the external radiation pressure of the Oph-S1 star acting on dust and gas in the densest part of the cloud, given by:
@@
where =1100 is the bolometric luminosity of the Oph-S1 star adopted from Lada & Wilking (1984) and 0.06 pc is the projected distance from the cloud to the Oph-S1 star. We obtained K cm-3, which is in agreement with the result calculated from the ionizing photons emitted by the Oph-S1 star (2 K cm-3, Pattle et al., 2015). Here, we omit the fact that the UV radiation from the Oph-S1 star can be absorbed by intervening dust in between the star and the Oph-A region. Thus, the actual acting on dust and gas in Oph-A is overestimated in the Eq. 16. To make a comparison with the external radiation pressure, we calculated the internal thermal pressure due to the hot gas in Oph-A, . We considered the hot gas within the densest part of Oph-A, where Stokes-0.4 Jy/arcsec2 at 154 m only, which corresponds to the area inside the third contour shown in bottom right panel of Fig. 5. Within this region, we adopted in the range of 104–106 cm-3 and the average temperature of the hot gas 20 K, which yields in the range of (0.2–20) K cm-3. With this assumption, we note that could be underestimated, since there is a gas heating process in the cloud dominated by the UV radiation from the Oph-S1 star and by the photoelectric emission from dust exposed to the UV radiation (i.e., ). The highest internal thermal pressure of 20106 K cm-3 is seen toward the densest part of the cloud, associated with the position of the SM1 core. This suggests that the internal thermal pressure is strong enough to help the material refrain from external pressure. We note that, however, in the vicinity region away from the center (e.g., the tails of the ridge shown in Fig. 12), the internal thermal pressure is weaker, at least one order of magnitude less than the external one. In this case, the radiation pressure from the Oph-S1 star on the east side is dominant and able to push away all the matter into the surrounding environment toward the west side. However, it is clear that the material in the low-density region still survives, despite owning a low internal thermal pressure (see Fig. 12). This indicates that there must be another mechanism at play, for example, the presence of strong B-fields of a few mG in the cloud. We roughly estimated the magnetic pressure () to be 107–108 K cm-3 for B-field strengths between 1–2 mG. This indicates that the magnetic pressure is at least two orders of magnitude higher than the external pressure of 106 K cm-3, sufficient to support the matter against the external pressure along with the gas thermal pressure.
@@
We further investigated the interplay between B-fields and radiation directions on the evolution of the cloud. We determined the main radiation direction from the Oph-S1 star toward the densest region associated with the position of the starless core SM1 as the sky-projected direction from the Oph-S1 star and SM1 core. We found that the main radiation direction is 70°, measured east of north. The main B-field orientation toward the densest region of Oph-A of 74° is almost parallel to the radiation field. The strong radiation pressure from the Oph-S1 star pushes all the material away and creates a dense shell toward the west side. To prevent dust and gas compression in the dense region, the B-fields exert a magnetic pressure perpendicular to the B-field orientation, seen perpendicular to the direction of the radiation pressure. This suggests that the magnetic pressure does not help the matter resist the external pressure, but tends to stretch the structure out along the northwest-southeast direction, resulting in an elongated shape of the dense shell. In the west side, the external thermal pressure is blocked by the gas thermal pressure due to the high-density material, resulting in the ridge shape in the cloud with very high densities. These results are consistent with results from 3D MHD simulation, which predicted that if the B-field orientation is parallel or along with the radiation direction, the cloud will tend to have a flattened shape Henney et al. (2009). However, in the northeast and southeast regions, the offset angles between the B-field vectors and radiation field are larger; namely, the B-field orientation becomes perpendicular to the radiation direction. In this case, the magnetic pressure and tension induced by B-fields run parallel with the radiation pressure, suggesting magnetic pressure can support the matter to confront the radiative compression and prevent the dust and gas from being swept away. This results in bending the material structure and forming a dense ridge with an umbrella-like shape. Interestingly, in the case study of Oph-A the ionization source is very close to the cloud (¡0.1 pc) and the offset position angle between the radiation and B-fields is not homogeneous; instead, it varies from parallel to orthogonal. This allows us to closely examine how the interplay between the B-fields and radiation fields could influence the material structure of the cloud.
@@ 5 Summary
In this paper, we present our analysis of the Ophiuchus dark cloud complex, using the SOFIA/HAWC polarimetric observations toward the Oph-A molecular cloud in the L1688 active star-forming region. Our main results are as follows:
@@
Using the dust polarization maps at 89 and 154 m, we derived the map of B-field morphology in Oph-A. We found that B-fields are generally well-ordered and primarily perpendicular to the ridge of the cloud toward the higher density regime. We estimated in the entire region of Oph-A using the DCF method, ranging from 0.2 to 2.3 mG at 89 and to 2.5 mG at 154 m. We found that is prominent toward the densest region of the cloud associated with the starless core SM1 and decreases toward the outskirts. This is the first time we have successfully estimated the distribution of the B-field strengths in Oph-A, which is essential to understanding how B-fields vary in the molecular cloud globally. Our derived B-field morphology and strengths well agree with previous measurements using the sub-mm dust polarization (Kwon et al., 2018). We also estimated the strength using more advanced techniques, including the ADF, ST, and DMA, toward two subregions associated with the densest part of the cloud. While the strengths estimated with the ADF method were in agreement with those from DCF, the ST and DMA methods provided lower values of 4.5 and 2-3 times, respectively.
@@
We studied the role of B-fields relative to gravity and turbulence by calculating the maps of the mass-to-flux ratio, Alfvénic Mach number, and plasma parameter toward Oph-A, using the inferred from SOFIA/HAWC data at 154 m. We found that toward the central dense parts the cloud is magnetically sub-critical and sub-Alfvénic where B-fields dominate over the gravity and turbulence. We obtained a result of in Oph-A, suggesting the cloud is supported by the strong magnetic pressure and much greater than the thermal gas pressure. The lower strengths obtained with the ST and DMA methods do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation. Our virial analysis in the entire cloud showed the virial parameter 15, suggesting that the cloud is gravitationally unbound and able to against the self-gravitational collapse. The cloud may be in the magneto hydrostatic equilibrium and may undergo gravitational collapse due to turbulent fluctuations on smaller scales. We should notice that our rough estimation of is a global parameter, applied for the entire cloud. In contrast, Oph-A is known as one of the densest parts of the active star-forming region L1688 hosting a few pre- and proto-stellar cores and embedded protostars. Higher spatial resolution observations toward the smaller scale region in the cloud will be necessary to investigate the dynamical state of the cloud in greater detail.
@@
We examined the role of B-fields in supporting the cloud against the radiation pressure from the nearby high-mass star Oph-S1. We found that at the center part of the cloud, the magnetic pressure indeed does not play a role in helping the material to avoid the onset of radiation pressure. However, toward the northeast and southeast regions of the cloud, the magnetic pressure positively supports the cloud to prevent both the effects of radiative pressure and gas thermal pressure.